OpenStudy (anonymous):

Identify which equation is a quadratic equation. y = 3x + 4 y + x2 = (x − 2)(x + 7) y = −4x(x − 4) + 2x y + 3 = (x + 12)(x2 + 2)

OpenStudy (anonymous):

quadratic is a equation in which highest power of x is 2 so simplify each and see which equation has highest power of x 2

OpenStudy (anonymous):

this is like the twelfth time i have seen this exact same rather silly problem where does it come from?

OpenStudy (anonymous):

if you multiply this \[y = −4x(x − 4) + 2x\] out, you will get a quadratic

OpenStudy (anonymous):

this one \(y = 3x + 4\) is linear,there is no square

OpenStudy (anonymous):

Okay, thank you! :)

OpenStudy (anonymous):

this one \(y + x^2 = (x − 2)(x + 7) \) looks like it could be a quadratic, but is isn't because the \(x^2\) term on both sides will cancel

OpenStudy (anonymous):

this one \(y + 3 = (x + 12)(x^2 + 2) \) if you multiply out you will get an \(x^3\) term, so not quadratic

OpenStudy (anonymous):

i'd still like to know where this silly question comes from and why everyone seems to have it